To what temperature must 30.0dm3 of helium at 25°C be cooled at 1 atm pressure for its volume to
be reduced to 1.0dm3 at same pressure?
1
Expert's answer
2020-06-09T07:18:40-0400

Keeping the pressure constant,

"V\propto T" or,

"\frac{V_1}{T_1}=\frac{V_2}{T_2}" And "V_1=30 \ dm^3,T_1=25\degree C=298 K"

"V_2=1 \ dm^3,T_2=?"

On substituting values,"\frac{30}{1}=\frac{298}{T_2}" "\implies" "T_2=9.93\ K=(9.93-273)\degree C=-263.07\degree C"